2. (6 pts) Based on data about the growth of a particular species of tree, the following logarithmic model was determined:
h(t) = 5.902 ln(t) + 6.137, where t = age of tree in years and h (t) = height of tree, in feet.
(Note that "ln" refers to the natural log function) (explanation optional)
Using the model,

(a)At age 4 years, how tall is this type of tree, to the nearest tenth of a foot?




(b) At age 13 years, how tall is this type of tree, to the nearest tenth of a foot?


Answer by Theo(13342)

the formula is h(t) = 5.902 * ln(t) + 6.127.
t = age of tree in years.
h(t) = height of tree, in feet).

at age 4, the formula becomesw h(4) = 5.902 * ln(4) + 6.127 = 14.30890932 feet. = 14 feet.

at age 13, the formula becomes h(13) = 5.902 * ln(13) + 6.127 = 21.26533111 feet = 21 feet.

ln is the natural log function which is the invetrse of the exponential function.

by definition, ln(x) = y if and only if y = e^x.

for example, if x = 5, then e^5 = 148.4131591 and ln(148.4131592) = 5.

you can use your calculator to confirm.